OpenStudy (anonymous):

What is the focus of the parabola y = 1/8 (x – 2)(^2) – 4

OpenStudy (anonymous):

i got this

OpenStudy (anonymous):

We want it in the form: \[ \large 4p(y-k)=(x-h)^2 \]

OpenStudy (anonymous):

wrong thing y=1/8(x-2)^2 -4 +4 +4 y+4 =1/8(x-2)^2 *8 *8 8(y+4) = (x-2)^2 (4*2) (y+4) = (x-2)^2

OpenStudy (anonymous):

The vertex is \((h, k)\) and focus is \((h, k+p)\)

OpenStudy (anonymous):

So that would make the focus (2, -2)

OpenStudy (anonymous):

And directrix is \(y = k-p\).

OpenStudy (anonymous):

u get it

OpenStudy (anonymous):

@FlyinSolo_424

OpenStudy (anonymous):

thank you I get it

OpenStudy (anonymous):

needed to freshen up